OpenStudy (anonymous):

Can anyone help me with this? I will fan and medal! Solve the quadratic equation by the square root property. (2x + 5)2 = 49 A. {-6, 1} B. {0, 1} C. {-27, 27} D. {1, 6}

OpenStudy (campbell_st):

take the square root of both sides of the equation \[(x +a)^2 = b~~~then~~~~ x + a ~= ~\pm \sqrt{b}\] then subtract 5 from both sides

OpenStudy (campbell_st):

lastly, you will need to simpify the answer...

OpenStudy (campbell_st):

opps forgot to say... divide the final answers by 2

OpenStudy (anonymous):

Thank you! @campbell_st I chose C but it was wrong.. help?

OpenStudy (anonymous):

@AMYCARTER

OpenStudy (anonymous):

@playgames1122

OpenStudy (playgames1122):

im going with C

OpenStudy (anonymous):

I chose c, but it was wrong? help? @playgames1122

OpenStudy (playgames1122):

ask my sister @ChicagoHusky

OpenStudy (anonymous):

@ChicagoHusky can you help?

OpenStudy (anonymous):

The answer is A

OpenStudy (anonymous):

Thank you! @ChicagoHusky